OpenStudy (anonymous):

Find the inverse f(x)=3√(x-7)-9

OpenStudy (ranga):

Is -9 inside the square root or outside?

OpenStudy (anonymous):

sorry its x/7 and to be clear is not the square root its the square root with the 3 in it

OpenStudy (anonymous):

outside the square root

OpenStudy (ranga):

\[\Large f(x) = \sqrt[3]{\frac{ x }{ 7 }} - 9\] Is that the function?

OpenStudy (anonymous):

yes thats it

OpenStudy (ranga):

Okay. f(x) is same as y. So put y in the place of f(x) \[\Large y = \sqrt[3]{\frac{ x }{ 7 }} - 9\]

OpenStudy (ranga):

To find the inverse, first interchange x and y. That is, replace y with x and x with y.\[\Large x = \sqrt[3]{\frac{ y }{ 7 }} - 9\]Now solve for y. Can you do that?

OpenStudy (ranga):

Isolate y by adding 9 to both sides. Then cube both sides. The cube and the cube root will cancel each other on the right. Tell me what you get.

OpenStudy (anonymous):

cuberoot of (x+9)= y/7 right

OpenStudy (ranga):

Not cuberoot but CUBE. We are cubing both sides. The left side will be (x+9)^3 and the right side will be y/7

OpenStudy (anonymous):

ok so (x+9)^3=y/7 if im still trying to isolate y do i divide everything by 7?

OpenStudy (ranga):

Multiply both sides by 7

OpenStudy (ranga):

What do you get?

OpenStudy (anonymous):

(7x+63)^3=y ?

OpenStudy (ranga):

No need to take the 7 inside. You can leave it as a factor outside: 7(x + 9)^3 = y Switch the right and left sides: y = 7(x + 9)^3 That is your inverse function.

OpenStudy (anonymous):

thank you for the help!

OpenStudy (ranga):

You are welcome.